# Eyal Michaeli

### v1

@@ 6.1 Assessing the Relevance of FID in Generative Data Augmentation
In Table˜6, we report the FID values, calculated using augmentations alongside their respective real datasets, as well as the corresponding accuracy achieved with each augmentation method. We observe that generative baselines such as Real Guidance and ALIA achieve lower FID scores, which suggest a higher similarity to the real data distribution. We suspect that this is the result of generating images that closely mimic the original dataset. In contrast, our method, SaSPA, is designed to create diverse augmentations that substantially differ from the real images, leading to higher FID scores. Despite these higher FID values, as shown in Table˜6, SaSPA demonstrates superior performance enhancements in accuracy across datasets. This highlights the importance of evaluating generative augmentation methods not only based on realism and similarity to real images, as measured by FID, but primarily on their actual impact on model performance. In the next section, we further provide an alternative metric for generative data augmentation.
@@ 6.2 Evaluating Augmentation Diversity with LPIPS
Table˜7 demonstrates that SaSPA achieves significantly higher LPIPS scores compared to Real Guidance (RG) and ALIA, indicating that SaSPA introduces much greater diversity in the generated augmentations. This substantial increase in diversity is crucial for enhancing model robustness and performance [31]. Qualitative evidence can be found in Figure˜7.
@@ 7 Conclusion
We propose SaSPA, a generative augmentation method specifically designed for FGVC. Our method generates diverse, class-consistent synthetic images through conditioning on edge maps and subject representation. SaSPA consistently outperforms both traditional and recent generative data augmentation methods. It demonstrates superior performance across multiple settings, including multiple setups of the challenging and less-explored full dataset training, as well as in scenarios of contextual bias and few-shot classification. Limitations and future directions are discussed in Appendix˜A.
@@ C.1 Comparing SaSPA with More Augmentation Methods
Experimental Setup. As noted in Section˜4.5, we use diff-mix training setup. This setup employs ResNet50 [20] with a resolution of $448^{2}$ and ViT-B/16 [14] with a resolution of $384^{2}$, both of which are higher than the $224^{2}$ resolution we use across the paper. We incorporate the integration of ControlNet and BLIP-diffusion for Cars and CUB datasets. We do not use BLIP-diffusion for the Aircraft dataset as it proved to be a better option, as evidenced in Table˜4. The accuracies of diff-mix and other methods, as reported in Table 1 of the diff-mix paper [60], establish the benchmarks for our comparative analysis.
Comparison Methods. The compared methods, implemented by diff-mix, include (1) Real-Filtering (RF) and (2) Real-Guidance (RG), both proposed by He et al. [21]. RG is described in Section˜E.3, which they implement with a lower translation strength ($s=0.1$). RF is a variation of Real-Guidance that generates images from scratch and filters out low-quality images by using CLIP [41] features from real samples to exclude synthetic images that resemble those from other classes. (3) DA-Fusion [54] solely fine-tunes the identifier using textual inversion [18] to personalize each sub-class and employs randomized strength strategy ($s\in\{0.25,0.5,0.75,1.0\}$), and non-generative augmentation methods (4) CutMix [66] and (5) Mixup [67].
@@
Results. Results show that SaSPA outperforms all methods across both architectures when evaluated on Aircraft and Cars, despite diff-mix using heavy fine-tuning. Continuing the discussion on CUB evaluation in Section˜4.5, CutMix outperforms all methods when using the ViT-B/16 architecture, while diff-mix leads on ResNet50. Notably, SaSPA outperforms all other generative baselines on CUB except diff-mix using both architectures.
@@ C.2 Effect of Traditional Augmentations with SaSPA
Across our experiments, we combined SaSPA with the best traditional augmentation method, as described in Section˜4.1. To test how SaSPA behaves without augmentation and whether it depends on the best traditional augmentation, we evaluated its interaction with CAL-Aug [44] (the default traditional augmentation used in CAL), the best traditional augmentation, and no traditional augmentation at all. Note that for 3 out of 5 datasets, CAL-Aug is the best traditional augmentation, so we did not provide separate results for SaSPA with the best traditional augmentation for these datasets (as they are the same). From the results in Table˜9, we observe that using no traditional augmentation significantly under-performs compared to using CAL-Aug or the best traditional augmentation. Additionally, CAL-Aug proved to be a robust choice, yielding similar accuracy across all datasets, with only a slight decrease in performance for CompCars.
@@ C.3 Robustness to Deeper Networks
Our primary experiments utilized ResNet50 as the backbone architecture. To further evaluate how deeper and more performant networks might benefit from our augmentation method, we analyzed results across three FGVC datasets as detailed in Table˜10. The performance of both ResNet50 and ResNet101 architectures is presented. Results indicate that deeper networks, such as ResNet101, benefit from our augmentation method. Together with our comparison with diff-mix [60] using their training setup, which also employs the ViT-V/16 architecture [15], this demonstrates that our method is robust across a variety of architectures and training setups.
@@ C.4 Effect of Scaling the Number of Augmentations (M)
Results presented in Figure˜6 demonstrate a consistent increase in validation accuracy for both datasets as the number of augmentations increases. Notably, the Cars dataset shows robust performance improvements, even surpassing the results achieved by adding 25% of real data when $M=4$, indicating the effectiveness of the SaSPA augmentations in this context. For the Aircraft dataset, the accuracy nearly reaches the levels achieved by adding 25% more real data when $M=5$.
@@ C.5 Evaluating Different Prompt Strategies
To assess the effectiveness of our proposed prompt generation, we evaluated various prompt strategies, and the results are detailed in Table˜11. To accelerate the experimentation process, these experiments were conducted using only ControlNet with SD XL Turbo on 50% of the data. Five main strategies were compared: (1) Captions: Direct use of captions as prompts, leveraging BLIP-2 [28] for captioning, as demonstrated to be effective in prior work [25]. (2) LE (Language Enhancement) [21] and (3) ALIA [16] are described in Section˜E.3. (4–5) Our Method with and without appending artistic styles. The artistic style augmentation involves appending half of the prompts with the phrases “, a painting of <artist>”, where <artist> refers to renowned artists such as van Gogh, Monet, or Picasso. This approach aims to diversify textures and colors, prompting an increase in the model’s robustness.
@@ C.6 Will More Prompts Improve Performance?
The results in Table˜12 show no significant difference when using 200 prompts, indicating that 100 prompts are sufficient.
@@ C.7 Investigating the Potential of Newer Base Models
Recently, text-to-image diffusion models have made incredible progress, particularly those based on Stable Diffusion (SD) [46]. Notable advancements include SD XL [40] and SD XL Turbo [51]. In this section, we aim to explore the compatibility of Edge Guidance with other base models. Unfortunately, as these models are relatively new, BLIP-diffusion [27] has not yet released versions built upon them, preventing us from utilizing subject representation. However, as shown in Table˜4, our full pipeline without subject representation still achieves impressive results. Additionally, Table˜4 indicates that when subject representation is not used, it is slightly beneficial for most datasets, except CUB, to append half the prompts with artistic styles. Therefore, we adopt this strategy. In Table˜13, we experiment with SD v1.5, SD XL, and SD XL Turbo. Note that ControlNet versions for SD XL and SD XL Turbo are still experimental, and will require reevaluation as the models mature.
The results indicate that integrating Edge Guidance generally has a positive impact across base models, except on the CUB dataset, which aligns with our earlier findings in Table˜4. Additionally, SD XL and SD XL Turbo typically outperform SD v1.5, suggesting that more advanced base models may lead to further improvements in performance.
@@ C.8 Does Stopping Augmentation at Early Epochs Help?
A common technique when training with synthetic data is to first train using the synthetic data, then fine-tune on the real data [56, 38]. Inspired by this, we investigate stopping the data augmentation at earlier training epochs. Results are presented in Table˜14. For these ablation experiments on evaluating early augmentation stoppage, we used 50% of the real Aircraft dataset for faster experimentation. We find no benefit from early augmentation stoppage. There is a downward trend in accuracy when stopping early, with the worst results at the earliest epoch stopped (20 out of 140). We believe that the high diversity introduced by our augmentations reduces over-fitting, mitigating the need for an explicit domain adaptation strategy. As a result, the model continues to benefit from the augmented data throughout the training process.
@@ C.9 Choice of Conditioning Type
Using Canny edge maps [6] as a condition has proven effective for generating images with high diversity and class fidelity. Here, we experiment with a different kind of edges: Holistically-Nested Edge Detection (HED) edges [63]. Canny edges are more focused on detecting the intensity gradients of the image, often capturing finer details, whereas HED edges provide a more structured representation by capturing object boundaries in a holistic manner. We experiment with both types in Table˜15, using the default generation parameters without using BLIP-diffusion, and on 50% of the data for faster experimentation. Using Canny edges resulted in slightly higher validation accuracy on the Aircraft dataset.
@@ Appendix D Dataset details
We provide the number of samples for each dataset split used in our experiments in Table˜16. Additionally, we include the number of images for each background class (sky, grass, road) used to create the contextually biased training set, as shown in Table˜17. We utilize the dataset test split for the reported test. For datasets lacking a validation split (Cars, CUB, CompCars), we generate one by using 33% of the training set. Note that in the diff-mix training setup (Section˜4.5), the training datasets for CUB and Cars consist of the original splits, as they did not use a separate validation split.
@@ E.1 Experimental Setup
Data Generation. All generative methods use the Diffusers library [57]. We employ BLIP-diffusion [27] and ControlNet [68] for SaSPA. Besides the prompt, an edge map for ControlNet, and a reference image for BLIP-diffusion as inputs for our generation, BLIP-diffusion requires source subject text and target subject text as inputs. We simply use the meta-class (e.g., “Airplane” for Aircraft dataset, “Bird” for CUB) of the dataset for both source and target subject texts. For all other diffusion-based augmentation methods, we use Stable Diffusion v1.5 [46]. For all diffusion-based models, including SaSPA, we use the DDIM sampler [53] with 30 inference steps and a guidance scale of 7.5. Images are resized to ensure the shortest side is 512 pixels before processing with Img2Img or ControlNet. We set the ControlNet conditioning scale to 0.75. For text-to-image generation, images are generated at a resolution of 512x512. We generate $M=2$ augmentations per original image for each experiment, and we use augmentation ratio $\alpha=0.4$ for all datasets except CUB [58], for which we use $\alpha=0.1$ as evidenced to be better in Figure˜5. We use $k=10$ in the top-k Confidence filtering. We use four NVIDIA GeForce RTX 3090 GPUs for image generation and training.
Training. We follow the implementation strategy outlined in the CAL study [44], tailored for FGVC. Our models, based on the ResNet50 [20] architecture, are fine-tuned for each dataset. Initialization uses PyTorch [39] pre-trained weights from ImageNet [12]. Optimization is performed with an SGD optimizer, with a momentum of 0.9 and a weight decay of 10-5, over 140 epochs. We adjust the learning rate and batch size during hyper-parameter tuning to achieve the highest validation accuracy. Training images are resized to 224x224 pixels. Results are averaged across three seeds. Specific values of hyper-parameters are in Table˜18.
@@ E.3 More details on Generative Baselines
Real-Guidance (RG): This method achieves impressive few-shot classification performance. For prompt generation, an off-the-shelf word-to-sentence T5 model, pre-trained on the “Colossal Clean Crawled Corpus” [42] and fine-tuned on the CommonGen dataset [29], is utilized to diversify language prompts. The model is used in order to generate a total of 200 prompts based on the meta-class. For image generation, SDEdit with a low translation strength ($s=0.15$) is used. Filtering is performed using CLIP filtering, which is described in Section˜3.4.
@@ F.2 Filtering Strategies
This section elaborates on our filtering mechanisms that remove lower-quality augmentations that do not correctly represent the sub-class or the meta-class. Additionally, we compare the effectiveness of alternative filtering methods in Table˜19.
Predictive Confidence Filtering utilizes the baseline model’s confidence to filter out augmentations whose true label does not rank within the model’s top-k predictions (k = 10). This baseline model is selected based on optimal performance outcomes from a hyperparameter sweep, as described in Section˜E.1. The choice of k can affect the results: using too low of a k can result in excessive filtering, limiting augmentations to those the baseline model already handles well, whereas too high of a k results in insufficient filtering, allowing low-quality augmentations to pass through. Therefore, we ablate on k as well to find the optimal value. We show a visualization of this filtering method in Figure˜8.
We also evaluate other filtering methods, including CLIP filtering [21], Semantic Filtering, and ALIA confidence filtering [16], as described in Section˜3.4 and Section˜E.3.
@@
Results from employing the various filters are presented in Table˜19. Note that for faster experimentation, we used 50% of the data (hence the low accuracy). Observations include: (1) CLIP filtering leads to poorer performance than using no filter, likely because CLIP struggles with fine-grained concepts such as specific car model tail lights. (2) Our confidence filtering method achieves the best results at $k=10$. (3) Combining our confidence filtering with semantic filtering surpasses all other methods.

### v2

@@ 6.1 Assessing the Relevance of FID in Generative Data Augmentation
In Table 6, we report the FID values, calculated using augmentations alongside their respective real datasets, as well as the corresponding accuracy achieved with each augmentation method. We observe that generative baselines such as Real Guidance and ALIA achieve lower FID scores, which suggest a higher similarity to the real data distribution. We suspect that this is the result of generating images that closely mimic the original dataset. In contrast, our method, SaSPA, is designed to create diverse augmentations that substantially differ from the real images, leading to higher FID scores. Despite these higher FID values, as shown in Table 6, SaSPA demonstrates superior performance enhancements in accuracy across datasets. This highlights the importance of evaluating generative augmentation methods not only based on realism and similarity to real images, as measured by FID, but primarily on their actual impact on model performance. In the next section, we further provide an alternative metric for generative data augmentation.
@@ 6.2 Evaluating Augmentation Diversity with LPIPS
Table 7 demonstrates that SaSPA achieves significantly higher LPIPS scores compared to Real Guidance (RG) and ALIA, indicating that SaSPA introduces much greater diversity in the generated augmentations. This substantial increase in diversity is crucial for enhancing model robustness and performance [31]. Qualitative evidence can be found in Figure 7.
@@ 7 Conclusion
We propose SaSPA, a generative augmentation method specifically designed for FGVC. Our method generates diverse, class-consistent synthetic images through conditioning on edge maps and subject representation. SaSPA consistently outperforms both traditional and recent generative data augmentation methods. It demonstrates superior performance across multiple settings, including multiple setups of the challenging and less-explored full dataset training, as well as in scenarios of contextual bias and few-shot classification. Limitations and future directions are discussed in Appendix A.
8 Acknowledgments
This work was supported in part by the Israel Science Foundation (grant No. 1574/21).
@@ C.1 Comparing SaSPA with More Augmentation Methods
Experimental Setup. As noted in Section 4.5, we use diff-mix training setup. This setup employs ResNet50 [20] with a resolution of $448^{2}$ and ViT-B/16 [14] with a resolution of $384^{2}$, both of which are higher than the $224^{2}$ resolution we use across the paper. We incorporate the integration of ControlNet and BLIP-diffusion for Cars and CUB datasets. We do not use BLIP-diffusion for the Aircraft dataset as it proved to be a better option, as evidenced in Table 4. The accuracies of diff-mix and other methods, as reported in Table 1 of the diff-mix paper [60], establish the benchmarks for our comparative analysis.
Comparison Methods. The compared methods, implemented by diff-mix, include (1) Real-Filtering (RF) and (2) Real-Guidance (RG), both proposed by He et al. [21]. RG is described in Section E.3, which they implement with a lower translation strength ($s=0.1$). RF is a variation of Real-Guidance that generates images from scratch and filters out low-quality images by using CLIP [41] features from real samples to exclude synthetic images that resemble those from other classes. (3) DA-Fusion [54] solely fine-tunes the identifier using textual inversion [18] to personalize each sub-class and employs randomized strength strategy ($s\in\{0.25,0.5,0.75,1.0\}$), and non-generative augmentation methods (4) CutMix [66] and (5) Mixup [67].
@@
Results. Results show that SaSPA outperforms all methods across both architectures when evaluated on Aircraft and Cars, despite diff-mix using heavy fine-tuning. Continuing the discussion on CUB evaluation in Section 4.5, CutMix outperforms all methods when using the ViT-B/16 architecture, while diff-mix leads on ResNet50. Notably, SaSPA outperforms all other generative baselines on CUB except diff-mix using both architectures.
@@ C.2 Effect of Traditional Augmentations with SaSPA
Across our experiments, we combined SaSPA with the best traditional augmentation method, as described in Section 4.1. To test how SaSPA behaves without augmentation and whether it depends on the best traditional augmentation, we evaluated its interaction with CAL-Aug [44] (the default traditional augmentation used in CAL), the best traditional augmentation, and no traditional augmentation at all. Note that for 3 out of 5 datasets, CAL-Aug is the best traditional augmentation, so we did not provide separate results for SaSPA with the best traditional augmentation for these datasets (as they are the same). From the results in Table 9, we observe that using no traditional augmentation significantly under-performs compared to using CAL-Aug or the best traditional augmentation. Additionally, CAL-Aug proved to be a robust choice, yielding similar accuracy across all datasets, with only a slight decrease in performance for CompCars.
@@ C.3 Robustness to Deeper Networks
Our primary experiments utilized ResNet50 as the backbone architecture. To further evaluate how deeper and more performant networks might benefit from our augmentation method, we analyzed results across three FGVC datasets as detailed in Table 10. The performance of both ResNet50 and ResNet101 architectures is presented. Results indicate that deeper networks, such as ResNet101, benefit from our augmentation method. Together with our comparison with diff-mix [60] using their training setup, which also employs the ViT-V/16 architecture [15], this demonstrates that our method is robust across a variety of architectures and training setups.
@@ C.4 Effect of Scaling the Number of Augmentations (M)
Results presented in Figure 6 demonstrate a consistent increase in validation accuracy for both datasets as the number of augmentations increases. Notably, the Cars dataset shows robust performance improvements, even surpassing the results achieved by adding 25% of real data when $M=4$, indicating the effectiveness of the SaSPA augmentations in this context. For the Aircraft dataset, the accuracy nearly reaches the levels achieved by adding 25% more real data when $M=5$.
@@ C.5 Evaluating Different Prompt Strategies
To assess the effectiveness of our proposed prompt generation, we evaluated various prompt strategies, and the results are detailed in Table 11. To accelerate the experimentation process, these experiments were conducted using only ControlNet with SD XL Turbo on 50% of the data. Five main strategies were compared: (1) Captions: Direct use of captions as prompts, leveraging BLIP-2 [28] for captioning, as demonstrated to be effective in prior work [25]. (2) LE (Language Enhancement) [21] and (3) ALIA [16] are described in Section E.3. (4–5) Our Method with and without appending artistic styles. The artistic style augmentation involves appending half of the prompts with the phrases “, a painting of <artist>”, where <artist> refers to renowned artists such as van Gogh, Monet, or Picasso. This approach aims to diversify textures and colors, prompting an increase in the model’s robustness.
@@ C.6 Will More Prompts Improve Performance?
The results in Table 12 show no significant difference when using 200 prompts, indicating that 100 prompts are sufficient.
@@ C.7 Investigating the Potential of Newer Base Models
Recently, text-to-image diffusion models have made incredible progress, particularly those based on Stable Diffusion (SD) [46]. Notable advancements include SD XL [40] and SD XL Turbo [51]. In this section, we aim to explore the compatibility of Edge Guidance with other base models. Unfortunately, as these models are relatively new, BLIP-diffusion [27] has not yet released versions built upon them, preventing us from utilizing subject representation. However, as shown in Table 4, our full pipeline without subject representation still achieves impressive results. Additionally, Table 4 indicates that when subject representation is not used, it is slightly beneficial for most datasets, except CUB, to append half the prompts with artistic styles. Therefore, we adopt this strategy. In Table 13, we experiment with SD v1.5, SD XL, and SD XL Turbo. Note that ControlNet versions for SD XL and SD XL Turbo are still experimental, and will require reevaluation as the models mature.
The results indicate that integrating Edge Guidance generally has a positive impact across base models, except on the CUB dataset, which aligns with our earlier findings in Table 4. Additionally, SD XL and SD XL Turbo typically outperform SD v1.5, suggesting that more advanced base models may lead to further improvements in performance.
@@ C.8 Does Stopping Augmentation at Early Epochs Help?
A common technique when training with synthetic data is to first train using the synthetic data, then fine-tune on the real data [56, 38]. Inspired by this, we investigate stopping the data augmentation at earlier training epochs. Results are presented in Table 14. For these ablation experiments on evaluating early augmentation stoppage, we used 50% of the real Aircraft dataset for faster experimentation. We find no benefit from early augmentation stoppage. There is a downward trend in accuracy when stopping early, with the worst results at the earliest epoch stopped (20 out of 140). We believe that the high diversity introduced by our augmentations reduces over-fitting, mitigating the need for an explicit domain adaptation strategy. As a result, the model continues to benefit from the augmented data throughout the training process.
@@ C.9 Choice of Conditioning Type
Using Canny edge maps [6] as a condition has proven effective for generating images with high diversity and class fidelity. Here, we experiment with a different kind of edges: Holistically-Nested Edge Detection (HED) edges [63]. Canny edges are more focused on detecting the intensity gradients of the image, often capturing finer details, whereas HED edges provide a more structured representation by capturing object boundaries in a holistic manner. We experiment with both types in Table 15, using the default generation parameters without using BLIP-diffusion, and on 50% of the data for faster experimentation. Using Canny edges resulted in slightly higher validation accuracy on the Aircraft dataset.
@@ Appendix D Dataset details
We provide the number of samples for each dataset split used in our experiments in Table 16. Additionally, we include the number of images for each background class (sky, grass, road) used to create the contextually biased training set, as shown in Table 17. We utilize the dataset test split for the reported test. For datasets lacking a validation split (Cars, CUB, CompCars), we generate one by using 33% of the training set. Note that in the diff-mix training setup (Section 4.5), the training datasets for CUB and Cars consist of the original splits, as they did not use a separate validation split.
@@ E.1 Experimental Setup
Data Generation. All generative methods use the Diffusers library [57]. We employ BLIP-diffusion [27] and ControlNet [68] for SaSPA. Besides the prompt, an edge map for ControlNet, and a reference image for BLIP-diffusion as inputs for our generation, BLIP-diffusion requires source subject text and target subject text as inputs. We simply use the meta-class (e.g., “Airplane” for Aircraft dataset, “Bird” for CUB) of the dataset for both source and target subject texts. For all other diffusion-based augmentation methods, we use Stable Diffusion v1.5 [46]. For all diffusion-based models, including SaSPA, we use the DDIM sampler [53] with 30 inference steps and a guidance scale of 7.5. Images are resized to ensure the shortest side is 512 pixels before processing with Img2Img or ControlNet. We set the ControlNet conditioning scale to 0.75. For text-to-image generation, images are generated at a resolution of 512x512. We generate $M=2$ augmentations per original image for each experiment, and we use augmentation ratio $\alpha=0.4$ for all datasets except CUB [58], for which we use $\alpha=0.1$ as evidenced to be better in Figure 5. We use $k=10$ in the top-k Confidence filtering. We use four NVIDIA GeForce RTX 3090 GPUs for image generation and training.
Training. We follow the implementation strategy outlined in the CAL study [44], tailored for FGVC. Our models, based on the ResNet50 [20] architecture, are fine-tuned for each dataset. Initialization uses PyTorch [39] pre-trained weights from ImageNet [12]. Optimization is performed with an SGD optimizer, with a momentum of 0.9 and a weight decay of 10-5, over 140 epochs. We adjust the learning rate and batch size during hyper-parameter tuning to achieve the highest validation accuracy. Training images are resized to 224x224 pixels. Results are averaged across three seeds. Specific values of hyper-parameters are in Table 18.
@@ E.3 More details on Generative Baselines
Real-Guidance (RG): This method achieves impressive few-shot classification performance. For prompt generation, an off-the-shelf word-to-sentence T5 model, pre-trained on the “Colossal Clean Crawled Corpus” [42] and fine-tuned on the CommonGen dataset [29], is utilized to diversify language prompts. The model is used in order to generate a total of 200 prompts based on the meta-class. For image generation, SDEdit with a low translation strength ($s=0.15$) is used. Filtering is performed using CLIP filtering, which is described in Section 3.4.
@@ F.2 Filtering Strategies
This section elaborates on our filtering mechanisms that remove lower-quality augmentations that do not correctly represent the sub-class or the meta-class. Additionally, we compare the effectiveness of alternative filtering methods in Table 19.
Predictive Confidence Filtering utilizes the baseline model’s confidence to filter out augmentations whose true label does not rank within the model’s top-k predictions (k = 10). This baseline model is selected based on optimal performance outcomes from a hyperparameter sweep, as described in Section E.1. The choice of k can affect the results: using too low of a k can result in excessive filtering, limiting augmentations to those the baseline model already handles well, whereas too high of a k results in insufficient filtering, allowing low-quality augmentations to pass through. Therefore, we ablate on k as well to find the optimal value. We show a visualization of this filtering method in Figure 8.
We also evaluate other filtering methods, including CLIP filtering [21], Semantic Filtering, and ALIA confidence filtering [16], as described in Section 3.4 and Section E.3.
@@
Results from employing the various filters are presented in Table 19. Note that for faster experimentation, we used 50% of the data (hence the low accuracy). Observations include: (1) CLIP filtering leads to poorer performance than using no filter, likely because CLIP struggles with fine-grained concepts such as specific car model tail lights. (2) Our confidence filtering method achieves the best results at $k=10$. (3) Combining our confidence filtering with semantic filtering surpasses all other methods.